OpenStudy (anonymous):

A garden table and a bench cost $1000 combined. The cost of the garden table is three times the cost of the bench. What is the cost of the bench?

OpenStudy (anonymous):

It's 1000 for both. If the table is three times the cost of the bench then: Table= 3/4 (three quarters) Bench= 1/4 (One quarter) Divide the 1000 by 4 to get one quarter, = 250 To get the price of the table multiply by three (750) The bench is one quarter so it's 250

OpenStudy (uri):

3x+x =1000 4x=1000 x=1000/4